OpenStudy (anonymous):

Find f+g, f-g, fg, and f/g. Determine the domain for each function. f(x)=sqrt(x) g(x)=x+15

OpenStudy (anonymous):

f+g = sqrt(x) + x + 15. domain: [0, inf ) range: [15, inf ) f-g = sqrt(x) - x -15 domain: [0, inf ) range: (-inf, -15]

OpenStudy (anonymous):

fg = sqrt(x)*(x+15) domain: [0, inf ) range: [0, inf ) f/g = sqrt(x)/(x+15) domain: [0, inf ) range: [0, sqrt(15)/30 )